OpenStudy (anonymous):

Calculate the exterior angle of a regualr polygon in which the interior angle is five times the exterior angle. Hence state the number of sides in the polygon. Please help =).

OpenStudy (anonymous):

The exterior angle is always 360. The interior angle can be stated as: 180 * (n-2) for a regular n sided pylogon. notice the triangle n = 3 => 180 * (3-2) = 180 Now you'll have to solve for n: (180 * (n-2) ) = 5 * 360

OpenStudy (anonymous):

Can you please explain step by step =/

OpenStudy (anonymous):

I think the exterior angle refers to the angle that is outside (opposite) the interior angle. Thus it should be 360-angle of interior angle. angle_exterior = 360deg - angle_interior

OpenStudy (anonymous):

Do you mean for the question, that an exterior angle is 5 times the interior angle?

OpenStudy (anonymous):

the interior angle is 5 times the exterior angle

OpenStudy (anonymous):

I'm think the exterior angle of a regular polygon is constant and 360 deg. Well looking at: 180 * (n-2) = 5* 360 This means that: The interior angle of a n sided polygon is 180 * (n-2) The the exterior angle times 5 equals the interior: 5 * 360 if you solve for n you'll get 180(n-2) = 5*360 n-2 = 5*360 / 180 = 10 n = 10 + 2 = 12 The polygon has 12 sides :)

OpenStudy (anonymous):

I'm not 100% certain now. I'm comparing the sum of the exterior and interior. I'm not sure if the question asks for a single interior angle and a single exterior.

OpenStudy (anonymous):

The question is asking for the exterior angle and the number of sides of the polygon

OpenStudy (anonymous):

Exterior angle: 30 deg Number of sides: 12

OpenStudy (anonymous):

Thank you very much! =)